-This is all the information provided to the question

When a solid is dissolved in water, the water is not an explicit reactant because the water is a solvent and mostly doesn't react. We can write the dissolution of LiCl(s) in water as
LiCl(s) \rightarrow Li^+(aq) + Cl^-(aq) .
We can also write H_2O over the arrow to indicate that the water is the solvent. Recall that the standard state for solutes in solution is 1M. Use the table to compute the ΔH° and ΔS° for dissolving LiCl(s) in water.

ΔH°____ kJ
ΔS°____ J/K
Answer #1

Liel(s) to, ut (aq) + cr (aq.) DHO = (APJ) L + $(DHg) (--(Atif) lici = -278.5 +(-1674) – (409). LAH = - 36.9 Kilmol 1 ASO2 su

NOTE : kindly note table has not been provided. I have done question using standard values.

if you have some other values kindly use that.